# PuzzleVQA: Diagnosing Multimodal Reasoning Challenges Through Abstract Patterns

### v1

@@ Abstract
Large multimodal models extend the impressive capabilities of large language models by integrating multimodal understanding abilities. However, it is not clear how they can emulate the general intelligence and reasoning ability of humans. As recognizing patterns and abstracting concepts are key to general intelligence, we introduce PuzzleVQA, a collection of puzzles based on abstract patterns. With this dataset, we evaluate large multimodal models with abstract patterns based on fundamental concepts, including colors, numbers, sizes, and shapes. Through our experiments on state-of-the-art large multimodal models, we find that they are not able to generalize well to simple abstract patterns. Notably, even GPT-4V cannot solve more than half of the puzzles. To diagnose the reasoning challenges in large multimodal models, we progressively guide the models with our ground truth reasoning explanations for visual perception, inductive reasoning, and deductive reasoning. Our systematic analysis finds that the main bottlenecks of GPT-4V are weaker visual perception and inductive reasoning abilities. Through this work, we hope to shed light on the limitations of large multimodal models and how they can better emulate human cognitive processes in the future [^1] Our data and code will be released publicly at https://github.com/declare-lab/LLM-PuzzleTest..
@@
[…]
@@ 1 Introduction
Through our investigation of leading large multimodal models, we find that existing models are not able to generalize well to simple abstract patterns. Notably, even GPT-4V is able to solve less than half of the puzzles on average. Our analysis reveals that its main bottlenecks are weaker visual perception and inductive reasoning abilities. Hence, our main contributions include:
@@
    We introduce PuzzleVQA, an automatically generated and diverse collection of abstract puzzles with reasoning explanations.
@@
    Our experiments show that even advanced large multimodal models do not generalize well to abstract patterns, and we identify their reasoning bottlenecks.
@@ 2 Background: Cognitive Theories
2.1 Cattell-Horn Theory
@@
2.2 Piaget’s Stages of Cognitive Development
This theory Piaget (1976) can provide a framework for progressing from basic sensory experiences to complex abstract reasoning and problem-solving. While we note that the large multimodal models do not develop in the same organic and experiential manner as humans, we are guided to explore how the models can emulate different stages of cognitive abilities. Concretely, through abstract patterns, we can evaluate how the models perceive multimodal information, reason inductively to extrapolate from observations to broader generalizations, and apply general principles to deduce the solution for specific problems.
@@ 3.2 Design Considerations
To avoid potentially noisy annotations, we use an automatic approach with multimodal templates to generate each puzzle. For instance, given a visual layout and pattern of a template, we can automatically populate the template with the appropriate objects that demonstrate the pattern.
@@
Figure 3: Taxonomy of abstract puzzles in PuzzleVQA with sample questions, based on fundamental concepts such as colors and size. To enhance diversity, we design both single-concept and dual-concept puzzles.
@@ 3.3 Puzzle Construction
To construct each abstract puzzle, we leverage multimodal templates based on fundamental concepts including numbers, colors, shapes, and size. Following the formulation and design considerations previously discussed, we first define the layout and abstract pattern for the puzzle. Each template can be randomly populated with the specific objects to represent the underlying pattern through demonstrations, forming a specific puzzle instance. For example, to construct a color-based puzzle instance shown in Figure 2, we focus on the concept of colors and define the layout as a hexagonal arrangement of six parts, with the abstract pattern that spatially opposite parts must have the same color. Thereafter, the template can be randomly populated to satisfy the pattern with colors from a predefined list of possible colors. Lastly, the query is constructed based on the fundamental concepts in the abstract pattern.
@@ 3.4 Multiple-Choice Format
While we use straightforward objects in our puzzles, there may be a degree of ambiguity in the answer regarding specific colors or sizes. Hence, we standardize the puzzle format as multiple-choice questions, where all questions are provided with four options, with the exception of three options for size (small, medium, and large). To generate the incorrect choices for each question, we use heuristics such as randomly sampling numbers within the same magnitude of the answer, and further details can be found in Appendix A.1. We use the standard accuracy metric for evaluation.
@@ 3.5 Dataset Analysis
To ensure that the dataset contains diverse abstract patterns, we provide a taxonomy of 10 puzzle categories based on fundamental concepts including numbers, colors, shapes, and size. As shown in Figure 3, there are four categories of single-concept patterns. To extend the depth of PuzzleVQA, we also include dual-concept patterns, which would require models to relate two concepts in order to solve the puzzle. Within each category, we design two multimodal templates that can each be used to generate many unique puzzle instances. The full list of puzzle templates and examples of more puzzle instances can be found in Appendix A.2. To maintain a reasonable dataset size for evaluating large multimodal models, we generate 100 unique puzzle instances from each template. Thus, there are 2000 test instances in PuzzleVQA in total.
@@
Table 1: Evaluation results for single-concept abstract patterns.
@@
Table 2: Evaluation results for dual-concept abstract patterns.
3.6 Implementation Details
We utilize Python code along with packages like Pillow [^2] https://pypi.org/project/pillow/ to automatically generate puzzles. Leveraging these tools, we are able to create many different unique puzzle images and text questions for each given puzzle type by augmenting the base template and objects in the image. We will release the dataset publicly with a permissive license such as MIT license.
@@ 4.1 Inference Pipeline
To elicit reasoning steps from large multimodal models, we leverage zero-shot chain of thought (CoT) prompting (Kojima et al., 2022) with a prompt similar to ‘‘Let’s think step by step’’. As the model may not always follow the same multiple-choice answer format, we also employ a model-based answer extraction stage. Detailed examples of the prompts can be found in Appendix A.3. While the main focus of this work is to evaluate how models can generalize to abstract patterns without prior experience, we also include a one-shot setting in Appendix A.4 which demonstrates slight but inconsistent performance improvements.
@@
Numbers Colors Size Shapes Average
LLaVA-13B 26.0 29.0 29.0 26.0 27.5
   w/ Guided perception 28.5 31.5 45.0 29.5 33.6
   w/ Guided perception & induction 39.5 38.5 43.0 44.0 41.2
   w/ Guided perception & induction & deduction 31.5 79.0 56.0 67.0 58.4
Gemini Pro 32.5 32.0 33.5 40.0 34.5
   w/ Guided perception 51.5 34.5 43.0 45.0 43.5
   w/ Guided perception & induction 79.0 46.5 49.0 49.0 55.9
   w/ Guided perception & induction & deduction 90.5 85.0 82.0 66.0 80.9
Claude 3 Opus 47.0 32.5 33.5 44.5 39.4
   w/ Guided perception 62.5 86.5 76.5 65.5 72.8
   w/ Guided perception & induction 99.5 100.0 95.5 98.0 98.2
   w/ Guided perception & induction & deduction 99.5 99.5 99.5 95.0 98.4
GPT-4V 67.5 42.0 35.0 41.0 46.4
   w/ Guided perception 87.0 66.5 62.5 67.5 70.9
   w/ Guided perception & induction 100.0 98.0 95.5 96.0 97.4
   w/ Guided perception & induction & deduction 100.0 100.0 100.0 99.0 99.8
Table 3: Analysis results on single-concept abstract patterns. To understand multimodal reasoning bottlenecks, we progressively prompt the models with ground truth visual perception, inductive reasoning, and deductive reasoning.
@@ 4.2 Models
To investigate the reasoning ability of large multimodal models, we select the best-performing open-source and closed source models Yue et al. (2023):
@@
    LLaVA-13B Liu et al. (2023) is an large multimodal model which is based on the popular LLaMA Touvron et al. (2023a) foundation language model. We use the 1.5 version and default chat template.
@@
  3. 3.
@@
  4. 4.
    Claude 3 Opus [^3] https://www.anthropic.com/news/claude-3-family is the best model in the Claude 3 model family released by Anthropic. We use their publicly available API to query the “claude-3-opus-20240229” version of the model.
@@ 5 Results
In the evaluation on dual-concept puzzles, as shown in Table 2, GPT-4V stands out again with the highest average score of 45.5. It performed particularly well in categories such as "Colors & Numbers" and "Colors & Size" with a score of 56.0 and 55.0 respectively. Claude 3 Opus closely follows with an average of 43.7, showing strong performance in "Numbers & Size" with the highest score of 34.0. Interestingly, LLaVA-13B, despite its lower overall average of 31.1, scores the highest in the "Size & Shapes" category at 39.0. Gemini Pro, on the other hand, has a more balanced performance across categories but with a slightly lower overall average of 30.1. Overall, we find that models perform similarly on average for single-concept and dual-concept patterns, which suggests that they are able to relate multiple concepts such as colors and numbers together.
Figure 4: Case study on two sample predictions from GPT-4V. The example on the left shows visual perception shortcomings and the example on the right shows the faulty inductive reasoning of the model which identified a spurious pattern in the image.
5.1 Analysis
Given the lower performance of existing large multimodal models, this raises the natural question of why they are not able to reason well about abstract patterns. As shown in Figure 2, the stages of solving abstract puzzles can be generally decomposed into visual perception, inductive reasoning, and deductive reasoning. Hence, we analyze their reasoning bottlenecks by progressively providing the ground truth explanation in their prompts. Note that we omit the final answer in the deductive reasoning explanation to avoid making the question trivial, and the detailed prompts can be found in Appendix A.3. The result of our analysis is shown in 3. Overall, we observe that the models perform better when provided with ground truth explanations, which suggests that they are able to leverage the additional information.
@@ 5.1 Analysis
6 Case Study
@@ 7 Related Work
The recent surge in multimodal pretraining and fine-tuning approaches has led to the creation of various benchmarks. Benchmarks like VQA Antol et al. (2015) and OK-VQA Marino et al. (2019) aim at evaluating the basic perception and reasoning abilities of large multi-modal models. Meanwhile, benchmarks like MMMU Yue et al. (2023) offer an evaluation of LLMs’ proficiency across a wide range of tasks requiring expert-level multimodal understanding.
Our work concentrates on creating a benchmark specifically tailored to abstract reasoning, aiming to evaluate how models emulate cognitive abilities, particularly in tasks involving reasoning about abstract concepts and relationships. Existing benchmarks such as RAVEN Zhang et al. (2019), CLEVR Johnson et al. (2017), ConceptARC Moskvichev et al. (2023), and MiniSCAN Lake et al. (2019) also target abstract reasoning but focuses on different aspects. RAVEN presents visual matrices with abstract patterns, challenging models to identify patterns and complete missing elements. However, we note that RAVEN has a specific spatial layout and can be solved exactly with search algorithms. Compared to CLEVR which offers synthetic images and questions focusing on logical and commonsense reasoning, we focus on exploring how large multimodal models mimic cognitive processes. Compared to ConceptARC which focuses more on spatial concepts such as inside and outside, our dataset focuses on object interactions based on fundamental concepts such as colors and shapes. Lastly, MiniSCAN maps pseudo words to a sequence of color symbols, but is limited to color patterns.
What distinguishes our dataset, PuzzleVQA, from existing works is its systematic analysis of various reasoning skills, including perceptual, inductive, and deductive reasoning. Notably, our dataset not only provides ground truth answers but also includes image captions and pattern explanations that enable a systematic diagnosis of the reasoning skills that a model may lack.

### v3

@@ Abstract
Large multimodal models extend the impressive capabilities of large language models by integrating multimodal understanding abilities. However, it is not clear how they can emulate the general intelligence and reasoning ability of humans. As recognizing patterns and abstracting concepts are key to general intelligence, we introduce PuzzleVQA, a collection of 2000 puzzle instances based on abstract patterns. With this dataset, we evaluate large multimodal models with abstract patterns based on fundamental concepts, including colors, numbers, sizes, and shapes. Through our experiments on state-of-the-art large multimodal models, we find that they are not able to generalize well to simple abstract patterns. Notably, GPT-4V achieves a score of 46.4% on single-concept puzzles, which shows that state-of-the-art models struggle on our dataset. To diagnose the reasoning challenges in large multimodal models, we progressively guide the models with our ground truth reasoning explanations for visual perception, inductive reasoning, and deductive reasoning. Our systematic analysis finds that the main bottlenecks of GPT-4V are weaker visual perception and inductive reasoning abilities. Through this work, we hope to shed light on the limitations of large multimodal models and how they can better emulate human cognitive processes in the future [^1] Our data and code are released at https://github.com/declare-lab/LLM-PuzzleTest..
@@
Yew Ken Chia1,2thanks:   Yew Ken Chia is under the Joint Ph.D. Program between DAMO Academy and the Singapore University of Technology and Design., Vernon Toh Yan Han1, Deepanway Ghosal1, Lidong Bing2, Soujanya Poria1 1 Singapore University of Technology and Design, 2 DAMO Academy, Alibaba Group, Singapore
1 Introduction
@@
Through our investigation of leading large multimodal models, we find that existing models are not able to generalize well to simple abstract patterns. Notably, GPT-4V achieves a score of 46.4% on single-concept puzzles, which shows that state-of-the-art models struggle on our dataset. Our analysis reveals that its main bottlenecks are weaker visual perception and inductive reasoning abilities. Hence, our main contributions include:
@@
    We introduce PuzzleVQA, an automatically generated and diverse dataset of 2000 multimodal samples with reasoning explanations.
@@
    Our experiments show that even advanced large multimodal models do not generalize well to abstract patterns, and we show how to identify their reasoning bottlenecks.
@@ 2 Background: Cognitive Theories
2.1 Fluid and Crystallized Intelligence
@@
2.2 Cognitive Development
Piaget’s Stages of Cognitive Development Piaget (1976) can provide a framework for progressing from basic sensory experiences to complex abstract reasoning and problem-solving. While we note that the large multimodal models do not develop in the same organic and experiential manner as humans, we are guided to explore how the models can emulate different stages of cognitive abilities. Concretely, through abstract patterns, we can evaluate how the models perceive multimodal information, reason inductively to extrapolate from observations to broader generalizations, and apply general principles to deduce the solution for specific problems.
@@ 3.2 Design Considerations
To avoid potentially noisy annotations, we use an automatic approach with multimodal templates to generate each puzzle. For instance, given a visual layout and pattern of a template, we can automatically populate the template with the appropriate objects that demonstrate the pattern. As each puzzle instance is created based on the specific rules in the template, we can ensure that they do not contain annotation mistakes.
@@
Figure 3: Taxonomy of abstract puzzles in PuzzleVQA with sample questions, based on fundamental concepts such as colors and size. To enhance diversity, we introduce both single-concept and dual-concept puzzles.
@@ 3.3 Puzzle Construction
To construct each abstract puzzle, we leverage multimodal templates based on fundamental concepts including numbers, colors, shapes, and size. Following the formulation and design considerations previously discussed, we first define the layout and abstract pattern for the puzzle. Each template can be randomly populated with the specific objects to represent the underlying pattern through demonstrations, forming a specific puzzle instance. For example, to construct a color-based puzzle instance shown in Figure 2, we focus on the concept of colors and define the layout as a hexagonal arrangement of six parts, with the abstract pattern that spatially opposite parts must have the same color. Thereafter, the template can be randomly populated to satisfy the pattern with colors from a predefined list of possible colors. Lastly, the query is constructed based on the fundamental concepts in the abstract pattern. To demonstrate our puzzle generation pipeline, we include a detailed implementation in Appendix A.4, based on the puzzle in Figure 1.
@@ 3.4 Multiple-Choice Format
While we use straightforward objects in our puzzles, there may be a degree of ambiguity in the answer regarding specific colors or sizes. Hence, we standardize the puzzle format as multiple-choice questions, where all questions are provided with four options, with the exception of three options for size (small, medium, and large). To generate the incorrect choices for each question, we use heuristics including randomly sampling numbers within the same magnitude of the answer, and further details can be found in the supplementary material. We use the standard accuracy metric for evaluation.
@@ 3.5 Dataset Analysis
To ensure that the dataset contains diverse abstract patterns, we provide a taxonomy of 10 puzzle categories based on fundamental concepts including numbers, colors, shapes, and size. As shown in Figure 3, there are four categories of single-concept patterns. To extend the depth of PuzzleVQA, we also include dual-concept patterns, which would require models to relate two concepts in order to solve the puzzle. Within each category, we design two multimodal templates that can each be used to generate many unique puzzle instances. The full list of puzzle templates and examples of more puzzle instances can be found in the supplementary material. To maintain a reasonable dataset size for evaluating large multimodal models, we generate 100 unique puzzle instances from each template. Thus, there are 2000 test instances in PuzzleVQA in total. We conducted an analysis in Appendix A.5 and found that the chosen dataset size is large enough to be relatively robust to experimental variance.
3.6 Implementation Details
We utilize Python code along with packages like Pillow [^2] https://pypi.org/project/pillow/ to automatically generate puzzles. Leveraging these tools, we are able to create many different unique puzzle images and text questions for each given puzzle type by augmenting the base template and objects in the image. Example code snippets to generate the puzzles are included in the supplementary material, and we plan to release the dataset publicly with a permissive license such as MIT license.
@@ 3.6 Implementation Details
Table 1: Accuracy of large multimodal models for single-concept abstract patterns in PuzzleVQA.
@@
Table 2: Accuracy of large multimodal models for dual-concept abstract patterns in PuzzleVQA.
@@ 4.1 Inference Pipeline
To elicit reasoning steps from large multimodal models, we leverage zero-shot chain of thought (CoT) prompting (Kojima et al., 2022) with a prompt similar to ‘‘Let’s think step by step’’. As the model may not always follow the same multiple-choice answer format, we also employ a model-based answer extraction stage. Detailed examples of the prompts can be found in the supplementary material. Please note that our main experimental setting used in Table 1 and 2 involves only the questions and images as multimodal inputs. On the other hand, we progressively provide additional ground-truth information such as image captions in Section 5.1 to diagnose the multimodal reasoning bottlenecks.
@@ 4.2 Models
To investigate the reasoning ability of large multimodal models, we select the best-performing open and closed-source models Yue et al. (2023):
@@
    Qwen-VL-Chat (7B) Bai et al. (2024) is an open-source large multimodal model designed to perceive and understand both texts and images. We use the version with open model weights and default chat template.
@@
    LLaVA-13B Liu et al. (2023) is an large multimodal model which is based on the popular LLaMA Touvron et al. (2023a) foundation language model. We use the model weights of the 1.5 version and default chat template.
  3. 3.
@@
  4. 4.
    Claude 3 Opus [^3] https://www.anthropic.com/news/claude-3-family is released by Anthropic and the most highest-performing multimodal model in their model family. We use their publicly available API to query the “claude-3-opus-20240229” version of the model.
  5. 5.
@@ 5 Results
In the evaluation on dual-concept puzzles, as shown in Table 2, GPT-4V stands out again with the highest average score of 45.5. It performed particularly well in categories such as "Colors & Numbers" and "Colors & Size" with a score of 56.0 and 55.0 respectively. Claude 3 Opus closely follows with an average of 43.7, showing strong performance in "Numbers & Size" with the highest score of 34.0. Interestingly, LLaVA-13B, despite its lower overall average of 31.1, scores the highest in the "Size & Shapes" category at 39.0. Gemini Pro, on the other hand, has a more balanced performance across categories but with a slightly lower overall average of 30.1. Overall, we find that models perform similarly on average for single-concept and dual-concept patterns, which indicates that they also struggle with puzzles that require reasoning about multiple abstract concepts.
LLaVA-13BGemini ProClaude OpusGPT-4V0050505050100100100100AccuracyOriginal prompt (Only question & image)w/ Guided perceptionw/ Guided perception, inductionw/ Guided perception, induction, deduction
Figure 4: Analysis on multimodal reasoning bottlenecks. We progressively prompt models with ground-truth explanations for visual perception, inductive reasoning, and deductive reasoning.
5.1 Analysis of Multimodal Reasoning Bottlenecks
Given the lower performance of existing large multimodal models, this raises the natural question of why they are not able to reason well about abstract patterns. As shown in Figure 2, the stages of solving abstract puzzles can be generally decomposed into visual perception, inductive reasoning, and deductive reasoning. Hence, we analyze their reasoning bottlenecks in Figure 4 by progressively providing the ground truth explanation in their prompts. Note that we omit the final answer in the deductive reasoning explanation to avoid making the question trivial, and the detailed prompts can be found in the supplementary material. Overall, we observe that the models perform better when provided with ground truth explanations, which suggests that they are able to leverage the additional information.
@@ 5.1 Analysis of Multimodal Reasoning Bottlenecks
0020202020404040406060606080808080100100100100Human BaselineGPT-4VClaude 3 OpusGemini Pro91.691.691.691.647.547.547.547.542.542.542.542.527.527.527.527.5Accuracy
Figure 5: Comparison between average human performance and large multimodal models on a subset of PuzzleVQA.
5.2 Comparison to Human Performance
To further shed light on how the large multimodal models compare to the reasoning ability of humans, we conducted a human baseline study involving 23 university students [^4] Note that the participants volunteered for the short study and we obtained prior permission from their instructor.. Participants were allotted 30 minutes to solve 40 puzzle instances sampled from our 20 puzzle categories, yielding an average human baseline score of 91.6%, as shown in Figure 5. Note that the 20-24 age group of the participants correspond to the formal operational stage of cognitive development, as discussed in Section 2. In contrast, the highest-performing model, GPT-4V scored 47.5% on the same set of puzzle samples, highlighting the specific bottlenecks causing models to fall short of human cognition: primarily in visual perception and inductive reasoning, as discussed in Section 5.1.
0123420202020303030304040404050505050Number of DemonstrationsAccuracyGPT-4VGemini ProLLaVA-13B
Figure 6: Analysis on the effect of few-shot demonstrations on model performance for single-concept puzzles.
5.3 Effect of Few-Shot Demonstrations
While we focus on the zero-shot setting to investigate how multimodal models handle novel reasoning challenges, we also explore how models may use knowledge and strategies from other puzzles to solve a new, specific puzzle. This is akin to analogical reasoning, which involves using experience from similar scenarios to make inferences about a novel situation Webb et al. (2022). Concretely, we run an analysis to study the effect of in-context learning Brown et al. (2020) with few-shot demonstrations of other puzzle instances when the model is tasked to solve a specific puzzle. The demonstrations consist of interleaved instances of multimodal inputs of each puzzle image and question, as well as the ground truth reasoning explanations. To ensure that the demonstrations are sufficiently diverse, we randomly select puzzles of different categories from the given puzzle. As shown in Figure 6, we find a general trend of increasing performance with respect to the number of demonstrations. Although there are some cases of lower performance for GPT-4V, we see that models generally achieve their best performance with the most number of demonstrations. This suggests that the models are indeed capable of analogical reasoning, and in-context learning may be a promising direction to enhance the abstract reasoning abilities of multimodal models in the future.
In addition, while not the main focus of this work, there may be other methods of improving model performance, including model training or different prompting methods. Hence, we have also included preliminary studies on fine-tuning with LLaVA-13B and comparison between chain-of-thought and direct prompting in the supplementary material. To consider the effect of other factors, the supplementary material further includes an alternative setting with text-only models given the ground-truth visual perception caption, and an analysis on the effect of evaluation dataset size.
@@ 6 Related Work
The recent surge in multimodal pretraining and fine-tuning approaches Liu et al. (2023); Bai et al. (2024) has led to the creation of various benchmarks. Benchmarks like VQA Antol et al. (2015) and OK-VQA Marino et al. (2019) aim at evaluating the basic perception and reasoning abilities of large multimodal models. Meanwhile, benchmarks like MMMU Yue et al. (2023) and ScienceQA Lu et al. (2022) offer an evaluation of LLMs’ proficiency across multiple disciplines requiring domain-specific knowledge and multimodal understanding.
To investigate the fundamental challenges in multimodal perception and reasoning, we deliberately focus on the abstract domain, aiming to assess how models emulate cognitive abilities, particularly involving reasoning about abstract concepts and relationships. However, we note that existing benchmarks have limitations which make them less suitable for studying large multimodal models. The RAVEN dataset Zhang et al. (2019) presents visual matrices with abstract patterns, challenging models to identify patterns and complete missing elements. However, we note that it has a specific spatial layout and can be solved exactly with search algorithms. Compared to CLEVR Johnson et al. (2017) which offers synthetic visual scenarios and questions focusing on logic and commonsense, we focus on exploring how large multimodal models perceive and reason about multimodal patterns, which is more closely related to fundamental cognitive processes in humans Mattson (2014). While ConceptARC Moskvichev et al. (2023) focuses on specific spatial concepts such as inside-outside and above-below, our dataset PuzzleVQA studies how visual objects interact and relate based on broader abstract concepts such as colors, shapes, numbers, and size. Lastly, the MiniSCAN dataset Lake et al. (2019) presents patterns that map special words to a sequence of color symbols, but is limited to color-based patterns.
What distinguishes our dataset, PuzzleVQA, from the existing works is its systematic analysis of multimodal reasoning through abstract patterns, including perceptual, inductive, and deductive reasoning. Compared to the previous datasets, our multimodal patterns encompass broad and fundamental abstract concepts such as numbers, colors, shapes, and size. Notably, our dataset not only provides ground truth answers but also includes image captions and pattern explanations that enable more detailed and systematic diagnosis of the reasoning bottlenecks for large multimodal models.